OpenStudy (thedj4jc):

Please help!! Will give medal and fan!! A grocer wants to make a 10-pound mixture of peanuts and cashews that he can sell for $4.75 per pound. If peanuts cost $4.00 per pound and cashews cost $6.50 per pound, how many pounds of each should he use? Let p = pounds of peanuts and let c = pounds of cashews. Write a system of equations that could be used to solve the problem.

OpenStudy (anonymous):

p+c=10 so 4p+6.5c=4.75(10)

OpenStudy (thedj4jc):

Okay, that makes sense. How do I find the number of pounds for p and c?